OpenStudy (anonymous):

How do i solve |2x+3|=9 when there is supposed to be two solutions?

OpenStudy (anonymous):

first consider: 2x +3 = 9 2x = 9-3 2x = 6 x = 3 then also consider: 2x + 3 = -9 2x = -12 x = -6

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

How about |x|=4?

OpenStudy (anonymous):

also consider two cases: x = 4 and x = -4

OpenStudy (anonymous):

in general, |x| = x, if x > 0 |x| = -x, if x < 0

OpenStudy (anonymous):

that's why you must consider these two cases, because you don't know what x is